### v4

@@ Abstract
Security issuance through intermediaries is subject to informational and agency-related frictions. However, separating their effects on securities has been difficult. We estimate those effects separately using SPAC data. To that end, we identify “premium” investors who produce value-relevant information. Their participation is associated with lower liquidation risk and higher returns. In contrast, “non-premium” investors engage only in quid pro quo arrangements. They receive high returns from an intermediary (quid) for participating in weaker future deals initiated by that intermediary (quo). Thus, quid pro quo is not pure agency cost; it includes transfers enabling more firms to go public.
@@ 1 Introduction
Firms issuing securities through intermediaries face two frictions. First, an informational gap between them (issuers) and investors may prevent securities from being sold at full value. Second, there may be an agency problem if intermediaries do not maximize sale proceeds but instead engage in quid pro quo arrangements with investors. Specifically, intermediaries may favor certain buyers (investors) at the expense of the issuers, expecting their favoritism to be returned through profitable future transactions.
Ritter and Welch (2002) point out that both frictions coexist and emphasize the need and difficulty in separating those frictions. In a survey paper, Lowry et al. (2017) similarly pose this separation as an unresolved question, using IPO underpricing as an example:
@@
We revisit this problem and separately estimate the effects of these frictions in the context of special purpose acquisition companies (SPACs)–publicly traded shell companies formed by sponsors to acquire or merge with private targets to take them public.
Like underwriters in a traditional IPO, sponsors are intermediaries between a target firm wanting to go public and investors. In both cases, an intermediary negotiates with a private company to determine the offer price, number of shares sold, and allocation across investors. Likewise, they profit if and only if the merger is successful.
In our setting, we distinguish between two types of investors: “premium” and “non-premium” (defined shortly below). We find evidence consistent with premium investors acting in the information production role. Their certification is associated with successful mergers and higher announcement returns. Non-premium investors engage only in quid pro quo arrangements. Specifically, when non-premium investors are allocated shares from a “hot” deal, they participate in a weaker deal initiated by the same sponsor in the future. Thus, interestingly, this quid pro quo is not all value-destroying. It partly enables a wider range of issuers to access public markets.
@@
We can isolate the effects of informational frictions separately from the agency friction because the SPAC setting offers several research design advantages over the traditional IPO. First, a SPAC divides the process of going public into distinct phases. In the initial stage, the sponsor raises capital from investors. A SPAC starts as a newly formed shell company without an operating business, so there is limited scope for information asymmetry (about underlying business) compared to a traditional IPO that starts with an operating business.111Consistent with this, in our sample, we find a small (1.3%) underpricing at SPAC IPO. The next phase consists of a “business combination” in which the SPAC merges with a private business and introduces new institutional investors known as “PIPE” (private investment in public equity), which affords us the ability to isolate the effects of PIPE investors.
@@
Second, legally, SPACs are publicly traded firms. Consequently, we observe the allocation to each PIPE investor and the effect of major events (e.g., business combination) on share prices. These data enable us to measure the return history for every sponsor-PIPE pair and create an in-sample measure of past relationships.222In traditional IPOs, allocations to institutional investors are unreported, and the quarterly 13F filings are insufficient to quantify favoritism if preferred clients “flip” shares in the secondary market. Furthermore, 13F filings do not distinguish shares allocated at the IPO from shares purchased in the secondary market. Third, the quid pro quo channel is easier to observe in SPACs than traditional IPOs. Unlike the intermediaries (investment banks) in traditional IPOs, SPAC sponsors typically do not operate in other lines of business (e.g., trading and wealth management). To the extent that sponsors earn quid pro quo from investors in future SPAC deals, we can observe that in our data.
Using data on all 1,072 SPACs from January 2010 to December 2021, we classify PIPE investors into premium and non-premium types based on participation frequency and investment size using k-means clustering. All else equal, frequent and large investors are more likely to produce value-relevant information than others (Benveniste and Spindt, 1989).
@@
A challenge in estimating the effects of premium investors is that the set of investors working with a sponsor is non-random. Better or more experienced sponsors may choose to work with better premium investors. If so, the estimated effects of premium investors could be biased because it would include the effect of assortative matching. So, we estimate fixed effects for each sponsor and investor following the methodology of Abowd, Kramarz and Margolis (1999) and use those fixed effects in our regression analysis. Our identifying assumption is that conditional on the fixed effects, the sponsor and premium investor matching is random.
@@
First, we find evidence consistent with premium investors offering value-relevant information and playing certification roles. Their participation is associated with significantly less share redemption by public investors and a large positive announcement effect on the stock price. Our estimates suggest that one standard deviation (32.85 pps) increase in premium investors’ participation is associated with a 5.91 pps decrease in the redemption rate (10.5% of the sample mean) and a 4.40 pps increase in the announcement-day return (117.7% of the sample mean).
@@
Next, we consider the agency friction. We note that sponsors profit an average of $126.9 million when the merger is successful; otherwise, they lose their initial investment, which averages $8.2 million. Therefore, the sponsors benefit from investors’ participation, which improves the chances of a successful merger. In turn, sponsors may have incentives to offer discounted shares to those investors in strong deals in exchange for an implicit promise to help them out with weaker deals in the future. We find evidence that only the non-premium investors engage in quid pro quo arrangements.
@@
In particular, we use the instrumental variables approach to document three ways quid pro quo manifests in the data. First, for each sponsor and non-premium investor pair, money left on the table (acronym, “MLOT”) exhibits mean-reversion, i.e., non-premium investors profit less in the current deal when they have high returns on past deals, and vice versa. In particular, one dollar above average MLOT in the last deal leads to a $2.72 below average MLOT from the current deal. We interpret this mean-reversion as the sponsor building a relationship with some non-premium investors and drawing it down when needed.
@@
Second, when the deal is weak or “cold,” i.e., when the offer share price at IPO is higher than the price on the closing day of the merger, the sponsor allocates shares disproportionately to those who earned strong returns from that sponsor in the past. Moreover, the sponsor allocates non-premium investors with a strong relationship more securities when the deals are weak than when they are strong.
@@
Third, a sponsor with strong relationships with non-premium investors is more likely to succeed. More precisely, if sponsors allocate one standard deviation more MLOT above average ($300 million) to non-premium investors, the liquidation probability of the current deal falls by 48.93 pps. As a reference, the average liquidation rate in our sample is 17.82%, so this decrease is economically meaningful. We also implement several robustness exercises, including placebo exercises, and find that our results are consistent with our hypotheses.
@@
Thus, we establish a new institutional feature in security issuance based on the relationship between the intermediary and institutional investors. In particular, we provide evidence that the quid pro quo arrangement enables cross-deal subsidies from high-quality target firms that will allow low-quality firms to go public. In contrast, the existing literature treats quid pro quo as a pure agency cost borne by issuers.
In summary, premium and non-premium investors play complementary roles and are paid accordingly. A premium investor, on average, is paid $6.96 million to produce information for the current target. A non-premium investor is paid $3.17 million to reduce liquidation risk in the future when the deal is weak, and premium investors are less likely to participate.
@@
Our findings about the differential effects of premium and non-premium investors complement the findings in Wang and Yung (2011) and Chemmanur, Hu and Huang (2010). Additionally, our finding that information production coexists with quid pro quo is reminiscent of Jenkinson, Jones and Suntheim (2018). Using data on the initial allocations of investment banks participating in traditional IPOs, they find evidence that investors are rewarded more when they submit price-limited bids and participate more actively in the meetings before or during bookbuilding. Both behaviors are consistent with an informational role. However, they also find that allocations are affected by how much revenue the investor generates for the bank, which is consistent with the quid pro quo argument. Our notion of quid pro quo differs slightly from the literature. Jenkinson, Jones and Suntheim (2018) document that investors build a relationship with a bank to get underpriced shares. We instead document that sponsors build relationships with non-premium PIPE investors to use them to complete future deals.
We also contribute to a nascent literature that seeks to understand the role and functioning of SPACs. For instance, Gahng, Ritter and Zhang (Forthcoming) study the cost and benefit of using SPACs to go public; Gofman and Yao (2022) study the conflict of interest and associated harm from directors serving on the boards of multiple SPACs; and Klausner, Ohlrogge and Ruan (2022) analyze the structure of SPACs and the associated costs.
@@ 2 A Primer on SPACs
During their IPOs, SPACs offer securities to public investors called units, typically priced at $10 per unit.333One exception is Periphas Capital Partnering Corporation, which priced its unit at $25. A unit comprises one Class A common stock share and a proportion (e.g., 1, 1/2, 1/3) of one redeemable warrant. Initially, these units trade as a whole. However, usually around the 52 calendar day following the IPO, owners have the option to either continue to hold units as a whole or trade shares and warrants separately.444After the separation, SPACs do not issue fractional warrants and only whole warrants trade. In other words, the number of warrants issued after separation equals the fraction rounded down to the nearest whole number. For example, if a unit contains one share and 1/3 of a warrant, then an investor with 1 unit receives no warrant, but an investor with 30 units receives ten warrants. Once the merger deal between the SPAC and a target is closed, a warrant gives an investor the right to purchase additional shares of common stock. The strike price is typically set well above the original SPAC price; a common choice is $11.50 per share. Any shares created in this manner automatically convert into ordinary shares after a merger.
@@
A SPAC generally has 24 months from its IPO to identify a target company, negotiate a deal, and complete the merger. Failure to complete the merger on time leads to its liquidation, and the SPAC must redeem all public shares and pay its investors with the funds held in the trust account, including any interest earned.
@@
Next, SPAC shareholders vote on the deal, and the transaction occurs only if a majority votes in favor. Shareholders can also redeem their shares, typically having at least 20 business days before the merger is complete. The merger agreement with the target company requires that the SPAC have a minimum net worth or a certain amount of cash (per the negotiation) as a closing condition. Sometimes, the funds raised in the SPAC IPO are sufficient to purchase the target the sponsor has located. More typically, the sponsor needs to raise additional funds. Because the SPAC is a publicly traded firm, these funds are called PIPEs.
@@
We categorize PIPE transactions based on the timing of their commitment: before, at, and after the business combination. In particular, the sponsor can invite PIPE investors before the SPAC IPO to commit to making certain investments in connection with the business combination. The PIPE transactions formed at the time of the IPO but before a target is identified are called forward purchase agreements; henceforth, they are PIPE (pre). The second type of PIPE transaction is announced along with the news about the business combination; henceforth, PIPE (at). The third type of PIPE transaction occurs after the target is announced but before the merger is completed; henceforth, PIPE (post).
A SPAC can have any combination of these three types of PIPEs. For example, Table 1 shows an example of SPAC, CF Capital Acquisition Corp V, which had all three PIPEs.
@@ 3 Testable Hypotheses
We develop testable hypotheses based on two competing paradigms – the information production and quid pro quo arrangements – that we assess in our data.
Information Production. Under the first paradigm, investors have or produce private information about the issuer and earn rents accordingly. This paradigm views a central goal of the going public process as extracting value-relevant information about the issuer and incorporating this information into the price.
Going public via SPACs involves investors in three stages (see Figure 1), and information production may have different implications at each step.
In particular, information production predicts zero underpricing during the SPAC IPO because a target has yet to be identified, and thus, there is no due diligence to perform. There is virtually no scope for investors to have private information; they exchange cash for a proportion of a trust fund of a fixed value.
In the business combination or de-SPAC phase, investors that participate in the target announcement (PIPE (at)) are the analogs to the buyers in the IPO bookbuilding models. Specifically, the sponsor finds a target and approaches PIPE (at) investors to solicit their interest. Then, the sponsor builds a book of PIPE demand, sets prices, and allocates shares.
However, these investors may differ in their abilities to produce information. For instance, Benveniste and Spindt (1989) suggest that premium investors who frequently access the market are likely more informed than others. In our context, such investors have private information about the quality of the issuing (target) firm and overall conditions affecting the market valuation. Conversely, non-premium investors are less informed and consequently have less impact on prices. Under these assumptions, the securities of the target firm may be underpriced to compensate premium investors for truthfully disclosing their private information about the target.
Furthermore, a premium PIPE (at) investor is incentivized to purchase more shares when its private information indicates the de-SPAC deal is profitable. Additionally, sponsors prioritize allocating to premium investors to encourage them to disclose their information. Consequently, a larger allocation to premium investors signals the good quality of the deal, which reduces redemption by public investors and leads to a positive market reaction. However, as non-premium PIPEs have no certification role, their participation should neither affect redemption nor lead to higher announcement-day returns.
@@
Under the information production paradigm, the following should hold:
@@
For a sponsor-premium PIPE (at) pair, MLOT is not mean-reverting.
Quid Pro Quo. We now discuss the implications of the quid pro quo paradigm. Under this paradigm, the seller of securities acts through a self-interested intermediary. Rather than selling securities at full price to premium investors, the intermediary gives discounts to some buyers, with an implicit understanding that this favor is rewarded through profits on future SPAC transactions.
The SPAC IPO presents the underwriter with such an opportunity. Underwriters likely deal with institutional investors more frequently (and in more varied capacities) than sponsors. As an intermediary, the underwriter may favor institutional investors (buyers) over sponsors (sellers), just as in a traditional IPO, resulting in SPAC underpricing.
The business combination phase presents another opportunity to observe favoritism, except the actors are swapped. The issuer is the private firm rather than the SPAC because the private firm, not the SPAC, issues new securities. The buyers are the PIPE investors who infuse new capital in exchange for securities. The intermediary is the sponsor, which negotiates terms and profits with PIPE investors only if the deal is completed.
Suppose that the underpricing at the business combination phase is not because of information production but because the sponsor allocates rents to PIPE (at) investors. Then, the sponsor should receive returns from these investors in the future. Then, including PIPE investors in a SPAC would not benefit the SPAC. Therefore, allocation to any PIPE (at) investors should not be associated with either redemption rates or announcement-day returns.
In any given deal, sponsors can favor particular PIPE investors and insist on favorable terms from the target firm. Favoritism may manifest differently at this stage than in traditional IPOs, where underwriters operate in many business lines, including trading and wealth management. An institutional investor receiving underpriced shares in a traditional IPO may repay the underwriter by providing business through other channels. In contrast, the sponsors operate only in the SPAC industry, and the primary objective of a SPAC sponsor is to close the existing fund and start future SPACs
We hypothesize that quid pro quo arrangements are employed only with non-premium PIPE investors. A sponsor showing favoritism to a particular non-premium PIPE investor expects reciprocity in future deals. Specifically, non-premium PIPE investors only participate in an underpriced deal (acting as a tax on the current private firm) in exchange for implied participation in a future weak deal. If a PIPE is involved only in “strong” deals, then the relationship is one-sided and involves “quid” with no “quo.”
While a long-term relationship (e.g., Levin, 2003; Halac, 2012) between a non-premium PIPE investor and a sponsor can provide necessary incentives to the investors to participate in weaker future deals, the observed interactions in any data are likely finite. Then, the question remains whether non-premium investors have incentives to participate in weaker future deals. In our context, it is appropriate to view non-premium investors as long-lived players who interact with sponsors who are short-lived players with perfect public monitoring. Such investors do not need to have repeated interactions with a particular sponsor to have the incentive to honor their commitment to participate in weak future deals (Fudenberg and Levine, 1994). If they renege to join in a weak deal led by a sponsor from whom they have received favorable returns, they risk losing future SPAC business.
Furthermore, under quid pro quo arrangements, the sponsor benefits from creating goodwill among investors, hoping to “draw down” on this goodwill in the future when the deals are weak. So, we hypothesize that a SPAC’s liquidation risk is negatively correlated with the amount of MLOT investors have received in the past from the sponsor of this SPAC.
@@
Under the quid pro quo paradigm, the following should hold:
@@
Allocations to non-premium PIPE (at) investors are uncorrelated with redemption and announcement-day returns.
@@
For a given sponsor -non-premium PIPE (at) pair, MLOT is mean-reverting.
@@
When a deal is weak, the sponsor is likelier to allocate shares to non-premium PIPE investors who earned large past MLOT.
@@
SPACs are less likely to liquidate when the sponsor has earned goodwill through high past MLOT. This connection occurs through only non-premium and not premium PIPEs.

### v5

@@ Abstract
Securities issuance through intermediaries is subject to agency problems and informational frictions. We examine these effects using SPAC data. We identify “premium” investors whose participation is linked to lower liquidation risk, higher returns, and lower redemption rates, consistent with both informational rents and agency frictions. In contrast, “non-premium” investors engage in non-agency quid pro quo relationships. Specifically, they receive high returns from an intermediary (quid) in exchange for a tacit agreement to participate in weaker future deals (quo). These relationships serve as insurance for issuers and intermediaries, enabling more issuers to access markets.
@@ 1 Introduction
One common view is that IPO issuers face an agency problem in which underwriters do not maximize sale proceeds but instead engage in quid pro quo arrangements with investors. Specifically, underwriters offer low-priced shares to favored investors, expecting to be repaid in other transactions (Ritter and Welch, 2002) benefiting the underwriter rather than the issuer. Drawing upon earlier work,111See, for example, Benveniste, Busaba and Wilhelm (2002) and Benveniste et al. (2003). Ljungqvist (2007) posits a second type of quid pro quo:
A second advantage of repeated interaction is that it allows underwriters to ‘bundle’ offerings across time. To ensure continued access to lucrative IPOs in the future, investors will occasionally buy poorly received IPOs as long as the loss they suffer in any given IPO does not exceed the present value of future rents they expect to derive from doing business with the underwriter.
Under the first view (“agency theory”), quid pro quo causes a deadweight loss in the securities market. It acts as a tax on issuers, decreasing their incentive to raise capital. By contrast, under the second view, quid pro quo arrangements represent a transfer from high-quality to low-quality issuers. For issuers uncertain about their valuation, this arrangement can ex-ante be helpful. It acts as insurance, allowing them to raise capital even when given a lukewarm market reception. Under this view (“insurance hypothesis”), quid pro quo plays a central and beneficial role in the intermediation process.
In this paper, we provide evidence consistent with the insurance hypothesis. In doing so, our paper is the first paper to show evidence of a positive role of quid pro quo arrangements in intermediated finance. Although this theory should apply to any setting of security issuance, in this paper, we focus on special purpose acquisition companies (SPACs), which are publicly traded shell companies formed by sponsors to merge with private targets, thereby taking them public. Unlike standard IPO data, SPAC data have unique features (that we make precise shortly below) that allow us to observe quid pro quo arrangements directly.
In our analysis, the sponsor (rather than a bank) is the intermediary. Like underwriters in a traditional IPO, sponsors serve as intermediaries between buyers (investors) and sellers (the issuer). In both cases, an intermediary negotiates with the issuer to determine the offer price and number of shares sold and helps determine the share allocation across investors. In both cases, the intermediary profits only if the deal is completed.
The identification challenge we face is that investors may perform dual roles–they provide information while also engaging in quid pro quo arrangements. For example, suppose a SPAC stock price rises upon the announcement of a particular investor’s involvement. This increase could be due to investor certification or the quid pro quo relationship, where the investor lacks private information but is allocated hot deals. The latter also leads to a price increase contemporaneously with the allocation, which is unrelated to certification.
We discuss these challenges in more detail later, but to briefly preview our results, we conclude that it is difficult to separate these two phenomena. This observation echoes the sentiment of Lowry et al. (2017) in their IPO survey paper, who lament the lack of progress on this separation, stating:
@@
By contrast, as we make clear in our hypothesis design section, the insurance hypothesis makes distinct predictions from the other theories, and our contribution is to isolate its effects in our data.
To this end, we distinguish between two types of investors: “premium” and “non-premium” (defined shortly below). We find evidence consistent with premium investors acting only in either the certification or agency role. Their presence is associated with higher announcement period returns, less redemption, and a higher likelihood of a successful SPAC merger. By contrast, non-premium investors engage only in quid pro quo arrangements as described by the insurance hypothesis. Specifically, we track the returns in each sponsor-investor pair, showing that within each pair, returns are mean-reverting. This pattern is consistent with the buildup and drawdown of relationship capital as per our insurance hypothesis. Second, we show that when a deal is weak, sponsors allocate higher shares to non-premium investors who earned higher returns in the past from the same sponsors than other non-premium investors who have yet to work with those sponsors. Third, the buildup of goodwill in this fashion serves as insurance by reducing the likelihood that a current SPAC is liquidated.
To summarize, the insurance hypothesis form of quid pro quo is evident in our data, and this role is played only by non-premium investors. Further, it is not destructive to issuers but enables a wider range of issuers to access public markets.
@@
As mentioned before, the SPAC setting offers several advantages in terms of research design over the traditional IPO. First, a SPAC divides the process of going public into distinct phases. In the initial stage, the sponsor raises capital from investors. A SPAC starts as a newly formed shell company without an operating business, so there is limited scope for information asymmetry (about the target firm) compared to a traditional IPO that starts with an operating business. The next phase consists of a “business combination” in which the SPAC merges with a private business and introduces new institutional investors known as “PIPE” (private investment in public equity), allowing us to isolate their effects on SPACs.
@@
Second, legally, SPACs are publicly traded firms. Consequently, we observe the allocation to each PIPE investor and the effect of major events (e.g., business combination) on share prices. These data enable us to measure the return history for every sponsor-PIPE pair and create an in-sample measure of past relationships.222In traditional IPOs, allocations to institutional investors are unreported, and the quarterly 13F filings are insufficient to quantify favoritism if preferred clients “flip” shares in the secondary market. Furthermore, 13F filings do not distinguish shares allocated at the IPO from shares purchased in the secondary market.
Our data include the universe of (1,072) SPACs publicly listed from January 1, 2010, to December 31, 2021. We follow these SPACs and record all major events and market returns until February 17, 2023. To categorize investor types, we follow the argument of Benveniste and Spindt (1989), who argue that all else equal, frequent and large investors are more likely to produce value-relevant information than others. To directly implement this notion, we classify PIPE investors into premium and non-premium types based on participation frequency and investment size using k-means clustering. We refer to such investors as premium investors and estimate their effects on SPAC outcomes.333We consider an alternative approach to assess if the insurance hypothesis is sensitive to how we define premium investors. For each PIPE investor, we searched and counted the articles published in The New York Times between January 2010 and December 2021 that featured the investor. We then define the investors with the top 5% counts as premium investors. This alternative definition also supports the insurance hypothesis.
@@
A challenge in estimating the effects of premium investors is that the set of investors working with a sponsor is non-random. Better or more experienced sponsors may choose to work with better premium investors. If so, the estimated effects of premium investors could be biased because of the assortative matching. To address this issue, we first estimate fixed effects for each sponsor and investor as proxies for their (unobserved) qualities following the methodology of Abowd, Kramarz and Margolis (1999) and then use those fixed effects in our regression analysis as controls. Our identification assumption is that conditional on those fixed effects, the sponsor and premium investor matching is random.
@@
We have several interesting results that shed light on the workings of the securities market. First, we find evidence consistent with premium investors offering value-relevant information and playing certification roles. Their participation is associated with significantly less share redemption by public investors and a large positive announcement effect on the stock price. In particular, one standard deviation (32.85 pps) increase in premium investors’ participation is associated with a 5.91 pps decrease in the redemption rate (10.5% of the sample mean) and a 4.40 pps increase in the announcement-day return (117.7% of the sample mean).
@@
Sponsors profit an average of $126.9 million when the merger is successful; otherwise, they lose their initial investment, which averages $8.2 million. Therefore, the sponsors benefit from investors’ participation, which improves the chances of a successful merger. In this sense, the quid pro quo arrangement suggested by Ljungqvist (2007) may benefit not only issuers in the form of insurance but also the sponsors. Thus, sponsors may have incentives to offer discounted shares to those investors in strong deals in exchange for an implicit promise to help them out with weaker deals in the future. We find evidence that only the non-premium investors engage in such quid pro quo arrangements.
@@
For our second set of findings, we use the instrumental variables approach to document three ways quid pro quo manifests in the data. First, for each sponsor and non-premium investor pair, money left on the table (“MLOT”) exhibits mean-reversion, i.e., non-premium investors profit less in the current deal when they have high returns on past deals, and vice versa. In particular, one dollar above average MLOT in the last deal leads to a $2.72 below average MLOT from the current deal. We interpret this mean-reversion as the sponsor building a relationship with some non-premium investors and drawing it down when needed.
@@
Second, when the deal is weak or “cold,” i.e., when the offer share price at IPO is higher than the price on the closing day of the merger, the sponsor allocates shares disproportionately to those who earned strong returns from that same sponsor in the past. Moreover, the sponsor allocates more securities to non-premium investors with whom it has a strong relationship when the deals are weak than when the deals are strong.
@@
Third, a sponsor with strong relationships with non-premium investors is more likely to succeed. More precisely, if sponsors allocate one standard deviation more MLOT above average ($300 million) to non-premium investors, the liquidation probability of the current deal decreases by 48.93 pps. As a reference, the average liquidation rate in our sample is 17.82%, so this decrease is economically meaningful. Using several robustness exercises, including placebo exercises, we verify that our results are consistent with how we define premium and non-premium investors.
@@
Thus, we establish a new institutional feature in security issuance based on the relationship between the intermediary and institutional investors. A premium investor is paid on average $6.96 million per deal, which, our estimates suggest, is a payment to produce information for the current target. A non-premium investor is paid, on average $3.17 million to reduce liquidation risk if, in the future, the current sponsor wants to merge with a weak target. Thereby, we provide evidence that the quid pro quo arrangement enables cross-deal subsidies from high-quality targets to low-quality targets that help the latter to go public. In contrast, the existing literature treats quid pro quo as a pure agency cost for the issuers.
@@
Our approach is partly motivated by the view of Benveniste, Busaba and Wilhelm (2002) that underwriters strategically “bundle” IPOs to resolve information asymmetries among companies with similar technologies. They suggest that “leaders” in this context can pave the way for "followers" through information production, leading to a standoff where no company wants to go public first. This issue is mitigated by transfers between leaders and followers, facilitating waves of IPOs. Benveniste et al. (2003) support this theory by showing that while initial returns and IPO volume are positively correlated overall, this correlation turns negative among offerings influenced by a common valuation factor. Unlike these papers, our observation is at the sponsor-investor pair level. Using information on investor allocations, we provide evidence of relationships between individual actors in the IPO market.
Our findings about the differential effects of premium and non-premium investors complement the findings in Wang and Yung (2011) and Chemmanur, Hu and Huang (2010). Additionally, our finding that information production coexists with quid pro quo is reminiscent of Jenkinson, Jones and Suntheim (2018). Using data on the initial allocations of investment banks participating in traditional IPOs, they find evidence that investors are rewarded more when they submit price-limited bids and participate more actively in the meetings before or during bookbuilding. Both behaviors are consistent with an informational role. However, they also find that allocations are affected by how much revenue the investor generates for the bank, which is consistent with the quid pro quo argument. They do not study the effect of quid pro quo or document any bundling as we consider here.
We contribute to the emerging literature on SPACs, examining their role and mechanics. Gahng, Ritter and Zhang (2024) explore the costs and benefits of SPACs for going public; Gofman and Yao (2022) investigate conflicts of interest and harm from directors on multiple SPAC boards; and Klausner, Ohlrogge and Ruan (2022) study SPAC structures and costs. Alti and Cohn (2023) addresses the adverse selection in choosing between traditional IPOs and SPACs, modeling how SPAC sponsors “cream skim” the best issues via costly search.
@@ 2 A Primer on SPACs
During their IPOs, SPACs offer securities to public investors called units, typically priced at $10 per unit.444One exception is Periphas Capital Partnering Corporation, which priced its unit at $25. A unit comprises one Class A common stock share and a proportion (e.g., 1, 1/2, 1/3) of one redeemable warrant. Initially, these units trade as a whole. However, usually around the 52nd calendar day following the IPO, owners have the option to either continue to hold units as a whole or trade shares and warrants separately.555After the separation, SPACs do not issue fractional warrants and only whole warrants trade. In other words, the number of warrants issued after separation equals the fraction rounded down to the nearest whole number. For example, if a unit contains one share and 1/3 of a warrant, then an investor with 1 unit receives no warrant, but an investor with 30 units receives ten warrants. Once the merger deal between the SPAC and a target is closed, a warrant gives an investor the right to purchase additional shares of common stock. The strike price is typically set well above the original SPAC price; a common choice is $11.50 per share. Any shares created in this manner automatically convert into ordinary shares after a merger.
@@
A SPAC generally has 24 months from its IPO to identify a target company, negotiate a deal, and complete the merger. Failure to complete the merger on time leads to its liquidation and the SPAC must redeem all public shares and pay its investors with the funds held in the trust account, including any interest earned.
@@
Next, SPAC shareholders vote on the deal, and the transaction occurs only if a majority votes in favor. Shareholders can also redeem their shares, typically having at least 20 business days before the merger is complete. The merger agreement with the target company requires that the SPAC have a minimum net worth or a certain amount of cash (per the negotiation) as a closing condition.
Typically, the purchase price of the target firm exceeds the funds held by the SPAC. In such cases, the SPAC raises additional funds from various private investors. Because the SPAC is a publicly traded firm, these funds are called PIPEs. More rarely (12% of our sample), the SPAC purchases the target without involving PIPEs.
Sometimes, the funds raised in the SPAC IPO are sufficient to purchase the target the sponsor has located. More typically, the sponsor needs to raise additional funds.
@@
We categorize PIPE transactions based on the timing of their commitment: before, at, and after the business combination. In particular, the sponsor can invite PIPE investors before the SPAC IPO to commit to making certain investments in connection with the business combination. The PIPE transactions formed at the time of the IPO but before a target is identified are called forward purchase agreements; henceforth, they are PIPE (pre). The second type of PIPE transaction is announced along with the news about the business combination; henceforth, PIPE (at). The third type of PIPE transaction occurs after the target is announced but before the merger is completed; henceforth, PIPE (post). A SPAC can have any combination of these three types of PIPEs. For example, Table 1 shows an example of SPAC, CF Capital Acquisition Corp V, which had all three PIPEs.
@@ 3 Testable Hypotheses
We develop testable hypotheses based on the following theories:
Certification: reputable PIPE investors have private information and certify deals.
Agency theory: Sponsors engage in quid pro quo arrangements with privileged PIPE investors at the expense of target firms.
Insurance hypothesis: Sponsors engage in quid pro quo arrangements with relationship PIPE investors in hot deals in order to complete future weaker deals.
We discuss the predictions of each paradigm in turn.
Certification Theory. Under the first paradigm, investors have or produce private information about the issuer and earn rents accordingly. This paradigm views a central goal of the going public process as extracting value-relevant information about the issuer.
Going public via SPACs involves investors in three stages (see Figure 1), and information production may have different implications at each step. In particular, we predict zero underpricing during the SPAC IPO because a target has yet to be identified, and thus, there is no due diligence to perform. There is virtually no scope for investors to have private information; they exchange cash for a proportion of a trust fund of a fixed value. In the business combination phase, investors participating in the target announcement (PIPE (at)) are the analogs to the buyers in IPO bookbuilding models. Specifically, the sponsor finds a target and approaches PIPE (at) investors to solicit interest. The sponsor builds a book of demand and–in conjunction with the selling firm–negotiates prices and allocations, considering any informational rents that need to be paid.
However, investors may differ in their abilities to produce information. For instance, Benveniste and Spindt (1989) suggest that premium investors who frequently access the market are likely more informed than others. In our context, such investors have private information about the quality of the issuing (target) firm and overall conditions affecting the market valuation. Conversely, non-premium investors are less informed and consequently have less impact on prices. Under these assumptions, the securities of the target firm may be underpriced to compensate premium investors for truthfully disclosing their private information about the target.
Furthermore, a premium PIPE (at) investor is incentivized to purchase more shares when its private information indicates the de-SPAC deal is profitable. Additionally, sponsors prioritize allocating to premium investors to encourage them to disclose their information. Consequently, a larger allocation to premium investors signals the good quality of the deal, which reduces redemption by public investors and leads to a positive market reaction. However, assuming non-premium PIPEs have no certification role, their participation affects neither redemption nor announcement-day returns.
@@
Under the certification theory, the following should hold:
@@
For a sponsor-PIPE (at) relationship pair, MLOT is not mean-reverting.
Agency Theory. To develop predictions mirroring those of Hypothesis 1, we concentrate on the business combination phase, during which the sponsor acts as an intermediary between the selling firm (target) and investors. Under this paradigm, investors lack private information. Instead, the sponsor acts as a self-interested intermediary, negotiating with the selling firm (target) to ensure that underpriced shares flow to privileged PIPE investors.
Consistent with traditional IPO literature, we assume that the intermediary need not grant all investors the same level of privilege. Rather, the intermediary might only have relationships with certain investors allocated hot deals, whereas other (non-favored) investors receive fairly priced offerings. We refer to favored PIPEs as "relationship" investors and let the data determine their identity.
@@
Under the agency theory, the following should hold:
@@
Allocations to relationship PIPE (at) investors are associated with lower redemption and positive announcement-day returns.
@@
For a sponsor-PIPE (at) relationship pair, MLOT is not mean-reverting.
Note that these predictions mirror Hypothesis 1, but the causality is reversed. Under Hypothesis 1, the SPAC stock price rises because the market believes that an informed investor has private information. Under Hypothesis 2, the announcement of relationship investors coincides with the announcement of an underpriced target in a non-causal manner. Privileged investors receive underpriced shares but do not influence stock prices.
As pointed out in the introduction, the fact that these two theories make similar predictions makes it challenging to distinguish between them. The announcement of the target’s identity coincides with the announcement of PIPE (at) investors, frustrating our ability to disentangle the effects of each announcement.
Insurance Hypothesis. Under this paradigm, sponsors hope to offer a sequence of SPACs but cannot guarantee that every target they find is high-quality. To provide themselves insurance, when they obtain hot deals, they favor particular PIPE investors from whom they expect to return the favor.666Note that the enforcement of the implicit relationship agreement does not require a long history between a sponsor and an investor. Indeed, reneging on even one deal destroys relationship capital, assuming participation in a deal is publicly observable; see Fudenberg and Levine (1994). As with the agency theory, we are agnostic about which PIPE investors serve this relationship role but instead let the data decide.
Specifically, the sponsor benefits from creating goodwill among relationship investors, hoping to “draw down” on this goodwill in the future when the deals are weak. Consequently, we hypothesize that a SPAC’s liquidation risk is negatively correlated with the amount of MLOT that relationship investors have received in the past from the sponsor of this SPAC.
Hypothesis 3.
Under the insurance hypothesis, the following should hold:
1.
Allocations to relationship investors are uncorrelated with redemption and announcement-day returns.
2.
For a given sponsor-PIPE relationship pair, MLOT is mean-reverting.
@@
When a deal is weak, the sponsor is more likely to allocate shares to relationship PIPE investors who earned large past MLOT.
@@
SPACs are less likely to liquidate when the sponsor has earned goodwill through high past MLOT.